b) How velocity of a particle can be defined in Quantum mechanics? Derive a relation for it. What information can be obtained from this velocity about momentum operator of the particle?


1
Expert's answer
2021-07-13T11:39:02-0400

Answer

Velocity operator can be define as momentum operator divided by mass of partical.

Momentum operator can be given as


"\hat{p}=\frac{\hbar}{i}\frac{\partial}{\partial x}"


And velocity operator

"\hat{v}=\frac{\hat{p}}{m}=" "\frac{\hbar}{im}\frac{\partial}{\partial x}"